OpenStudy (anonymous):

y=x+5 and y=-2x-1 the ordered pairs?

OpenStudy (anonymous):

You mean the point of intersection?

OpenStudy (anonymous):

yes , the points ?

OpenStudy (anonymous):

y = x + 5 -y = 2x + 1 3x + 6 = 0 3x = -6 x = -2 Then plug it back in. y = x + 5 y = -2 + 5 y = 3 Therefore, the point is (-2, 3)

OpenStudy (anonymous):

okay , thanks ill know for the next questionn how to do it .

OpenStudy (anonymous):

You're welcome